# SPICE: A Synergistic, Precise, Iterative, and Customizable Image Editing Workflow

### v1

@@ Abstract
Recent prompt-based image editing models have demonstrated impressive prompt-following capability at structural editing tasks. However, existing models still fail to perform local edits, follow detailed editing prompts, or maintain global image quality beyond a single editing step. To address these challenges, we introduce SPICE, a training-free workflow that accepts arbitrary resolutions and aspect ratios, accurately follows user requirements, and improves image quality consistently during more than 100 editing steps. By synergizing the strengths of a base diffusion model and a Canny edge ControlNet model, SPICE robustly handles free-form editing instructions from the user. SPICE outperforms state-of-the-art baselines on a challenging realistic image-editing dataset consisting of semantic editing (object addition, removal, replacement, and background change), stylistic editing (texture changes), and structural editing (action change) tasks. Not only does SPICE achieve the highest quantitative performance according to standard evaluation metrics, but it is also consistently preferred by users over existing image-editing methods. We release the workflow implementation for popular diffusion model Web UIs to support further research and artistic exploration [^1] https://github.com/kenantang/spice.
Image Editing, Training-Free, Diffusion Models
@@ 1 Introduction
Image editing is the task of changing the content of an image according to a user’s requirements. An example is adding an apple of a specific size at a specific location on an image (Figure 1) [^2] Figures in this paper are provided in high resolution. Readers are encouraged to zoom in to examine the details.. A powerful image editing tool is vital for many applications from creative design to scientific research. In artistic workflows, photographers modify or remove objects in an image to enhance storytelling and visual impact. In machine learning research, image editing has been used in video editing (Fan et al., 2025), data augmentation (Hirota et al., 2024), and benchmark construction (Wu et al., 2024).
@@
Figure 1: SPICE enables a user to edit the image exactly as they want, and image details improve instead of deteriorating after many editing steps. The first row shows the full image of a 3000$\times$2000 resolution. The second row shows a 900$\times$600 region enlarged for better visibility. In this example, a user uses 9 editing steps to perform various edits, including structure change, object removal, object addition, object replacement, text addition, color change, and detail fixes. Steps 7, 8, and 9 together fix the fridge structure. The labels above the first row are not the true editing instructions but their abbreviated version. For example, in the “Add a Word” column, the user wants to add the specific word “suspicious” to the white bowl. The prompt is “An open fridge with food in it. A bowl with a word ‘suspicious’ on it.” The result aligns with the user’s requirement.
@@
Existing vision language models have achieved initial success on the image editing task (Brooks et al., 2023; Zhang et al., 2024). Such models usually take in an original image and user’s editing prompt as the input and output the edited image. In addition, some methods use a binary mask to improve the editing results (Wang et al., 2023; Zhao et al., 2024). However, for advanced artistic or research purposes that require more than one editing step, existing methods are disqualified by the following limitations. First, pixels outside the mask often deteriorate after editing. Second, the user cannot specify the precise size and location of an object by the mask. Third, models struggle with unusual editing tasks, such as adding a backpack to a bench facing away from the viewer (Figure 3). Taken together, these limitations make it extremely difficult for users to refine an image across multiple edits, as image degradation inevitably accumulates during iterative editing.
@@
To overcome these limitations, we propose SPICE, a training-free workflow that can be easily adopted in existing text-to-image diffusion models. Our workflow follows a structured three-step process: (1) mask generation, (2) color and edge hint generation, and (3) two-stage denoising, as shown in Figure 2. First, in the mask generation step, the user provides a mask to define the editing region and the context size, localizing the modifications while using essential contextual information from the original image. By constraining the editing region to the user-provided mask, SPICE strictly preserves image quality and avoid deterioration outside the mask.
Second, during color and edge hint generation, SPICE allows the user to provide arbitrarily detailed or simplified image-space information to the model. This goes beyond conventional image editing methods (Brooks et al., 2023; Zhang et al., 2024) that rely solely on textual prompts which can hardly provide the precise size and location according to the user’s requirement. In contrast to existing methods, SPICE integrates structured visual guidance from a hinted image, offering a level of control over size and positioning that exceeds what textual prompts can achieve.
@@
Finally, SPICE uses a two-stage denoising process, where a Canny edge ControlNet model (Zhang et al., 2023) in the early denoising steps integrates image-space hints, and a base diffusion model refines the output in the later steps to generate diverse variations, as shown in Figure 2(c). By synergizing the complementary strengths of both models, the two-stage denoising process enables precise and customizable editing, in which the model faithfully follows the user’s requirements on the location, size, and other properties of the edited object. This not only enhances the quality of photo-realistic images in complex tasks but also allows AI-generated artwork to move beyond stereotypical and fundamental limitations, such as predominantly portraying a single character or repeatedly erring on details.
@@
By overcoming the limitations, SPICE consistently achieves success in the iterative editing tasks of more than 100 editing steps. The editing steps are a mixture of semantic and structural editing steps, where the image quality is consistently improved. For the first time, our workflow enables scaling test-time compute (Snell et al., 2024) in image generation, where the user can decide how each unit of additional compute time should contribute to the final image quality. This user-friendly design distinguishes our workflow from contemporary test-time scaling approaches that rely on automatic search algorithms and rewards (Ma et al., 2025; Singhal et al., 2025), over which the user has no granular control.
@@
Despite the strong capability, our workflow can be easily integrated into all popular diffusion model Web UIs, such as Stable Diffusion Web UI Automatic1111, ComfyUI, and Stable Diffusion Web UI Forge. Unlike existing tools such as ADetailer [^3] https://github.com/Bing-su/adetailer or Regional Prompter [^4] https://github.com/hako-mikan/sd-webui-regional-prompter, which offer an overwhelming number of hyperparameters that can be difficult to navigate, SPICE provides a smaller set of hyperparameters for ease-of-use. Furthermore, SPICE introduces minimal computational overhead in each editing step, making editing as efficient as generating an image from text with the same model and sampling hyperparameters. This extremely low overhead allows our workflow to run on consumer GPUs, unlocking high-quality and iterative image editing for more users. Combining strong capabilities with low implementation and computational cost, our workflow provides a powerful yet accessible image editing tool for researchers and non-researchers alike with huge convenience and flexibility.
@@
Figure 2: By sketching a binary mask with context dots and a color & edge hint, users can effortlessly achieve realistic edits with SPICE. Subfigure (a) shows the overview of our workflow, while Subfigures (b) and (c) show the internal steps. In this example, the user requires a sunhat to be added next to the woman. First, the user sketches both a mask with context dots and a hinted image containing color and edge hints. The mask is automatically blurred after being sketched. Then, during the two-stage denoising step, the Canny and base models perform the early and late denoising steps, respectively. See Figure 5 for more examples of masks and hints.
@@ 2 Methods
SPICE is based on inpainting (Lugmayr et al., 2022), an operation that uses a diffusion model to replace a masked region on an image, conditioned on a prompt that describes the new image. Note that this prompt differs from the editing prompt, which uses a verb (e.g., add or remove) to describe the editing operation (Zhang et al., 2024). For example, when the editing prompt is “add an apple in the fridge”, the description prompt will be “an apple in a fridge”. In this section, we introduce the three key steps of our workflow, namely mask generation, color and edge hint generation and two-stage denoising. An overview is shown in Figure 2.
@@ 2.1 Mask Generation
We denote the original image to be edited as $I_{T}\in[0,1]^{H\times W\times C}$, where $T$ is the index of the editing step, $H$ and $W$ are the image height and width, and $C=3$ represents the RGB color channels. Other than a prompt $p$, traditional inpainting methods [^5] https://github.com/AUTOMATIC1111/stable-diffusion-webui require the user to provide a binary mask $M\in\{0,1\}^{H\times W}$, where 1 indicates the region to be edited. Given the mask and the prompt, the inpainting operation uses the following steps to generate an edited image. First, a bounding box of the region to be edited is calculated, and the bounding box is extended either vertically or horizontally to ensure its aspect ratio matches a user-specified resolution supported by the diffusion model, such as 1216$\times$832 (Podell et al., 2024). The user-specified resolution is usually larger than the extended bounding box. Next, pixels on $I$ and $M$ within the extended bounding box are upsampled to the user-specified resolution. Then, the diffusion model generates a new image from latent noise. The generation process is conditioned on the description prompt and the upsampled pixels from $I$ and $M$. Finally, the output is downsampled to the resolution of the extended bounding box, and the inpainted region on $I_{T}$ is replaced by the downsampled output, resulting in $I_{T+1}$.
@@
To address these issues, we introduce context dots, a pair of dots at opposite corners of the desired bounding box. Context dots ensure that the extended bounding box includes sufficient context for generation. The user directly adds context dots to the original mask, resulting in a context mask $M_{\text{context}}\in\{0,1\}^{H\times W}$, as shown in Figure 2(b). This simple enhancement offers three key advantages: (1) the user can exclude image areas that interfere with the inpainting process, (2) the user can specify a resolution between that of the inpainted region and that of the full image, balancing local details and global context (Section 3.4), and (3) context dots minimally affect surrounding pixels, limiting changes to only the desired editing region.
@@
On inpainted images, there is usually an unwanted sharp boundary between the inpainted region and the remaining parts of the image. In more severe cases, an inpainted object can be incomplete. The reason is that even with a context, diffusion models are not robust enough to generate pixels that seamlessly blend with existing ones. To mitigate these artifacts, we adopt Differential Diffusion (Levin & Fried, 2023), a method that allows the diffusion model to be conditioned by continuous mask values in [0, 1] during generation. We provide the continuous values as a soft mask $M_{\text{soft}}\in[0,1]^{H\times W}$ by applying a simple Gaussian blur to $M_{\text{context}}$, as shown in Figure 2(b). This procedure, together with thresholds for blending the original and inpainted image, is named as Soft Inpainting in popular Web UIs. We use Soft Inpainting when it is available in an Web UI or implement our own simplified version (without thresholds) when it is not available.
@@ 2.2 Color and Edge Hint Generation
In many existing image editing systems (Croitoru et al., 2023; Yang et al., 2023; Huang et al., 2024), users typically specify desired content through a textual prompt. However, words alone cannot easily describe certain details, such as asymmetric apparel designs or intricate color patterns. Instead, these subtle details can be effectively conveyed with an additional visual hint in the image space. To this end, SPICE allows a user to first create a rough sketch or color layout using standard editing software (e.g., Krita or Adobe Photoshop). The resultant image, denoted $I_{\text{hinted}}\in[0,1]^{H\times W\times C}$, then replaces the original image as the inpainting input. To incorporate information from this hinted image, a denoising strength hyperparameter in [0, 1] specifies how much the original pixels on the image should be changed [^6] The hyperparameter does not work as a simple linear blending coefficient at the post-processing stage, and its exact implementation differs for various algorithms. To avoid ambiguity, we refrain from providing a general equation here. Interested readers can refer to the source code of popular Web UIs for the equations. . At a moderate denoising strength (e.g., 0.5 to 0.7), the inpainting model produces pixels that remain close to the user’s sketched hints—preserving intended colors, shapes, or patterns—while still diverging enough from the sketch to form realistic objects.
Note that the user can provide any form of color and edge hints, including sketches, reference images pasted in a collage style, or even another region of the original image (using the Clone Stamp Tool in PhotoShop). This flexibility is an improvement over existing methods, such as MagicQuill (Liu et al., 2024b) that uses downsampled 32$\times$32 color blocks for guidance and thus loses high-frequency details. Hence, SPICE allows a user to perform image editing from any point on the human-model collaboration spectrum. At one end, the user fully edit the image by drawing out every detail, without the help of diffusion models. At the other end, the user fully delegates the model to edit the image. Usually, the user can get decent editing results by staying on the end where the user input is minimal (Section 3.1). We will also discuss how a small set of hyperparameters enable the user to move freely along this spectrum (Section 3.4).
@@ 2.3 Two-Stage Denoising
Image editing methods (Lugmayr et al., 2022) typically rely on a single diffusion model to perform all denoising steps when generating an image from latent noise. However, different diffusion models have complementary strengths. On the one hand, a general-purpose text-to-image base model, such as Flux.1 [dev] [^7] https://huggingface.co/black-forest-labs/FLUX.1-dev, excels at generating rich variations in its output but can only be conditioned on textual prompts. On the other hand, Flux.1 [dev] Canny [^8] https://huggingface.co/black-forest-labs/FLUX.1-Canny-dev, a model derived from Flux.1 [dev] that contains a Canny edge ControlNet (i.e., a Canny model), can be conditioned on Canny edge information (Zhang et al., 2023) but sacrifices variability.
@@
Empirically, we find that for a wide range of image editing tasks, Canny Edge ControlNet models outperforms other ControlNet variants (e.g., depth or pose) in the first denoising stage. Furthermore, Canny models are more widely available than other ControlNet models. Therefore, we only use Canny models but not other ControlNet models.
@@ 3 Experiments
Original IP2P MagicBrush UltraEdit Ours Original IP2P MagicBrush UltraEdit Ours
@@
Figure 3: Our workflow outperforms baseline methods in 6 editing categories from EditEval. Each group of five images shows an example from an editing task. Each group starts from original image, followed by IP2P, MagicBrush, UltraEdit, and our results.
3.1 Benchmark Results
Evaluation Benchmark.
We use the second version of the EditEval (Huang et al., 2024) benchmark to evaluate our workflow and baseline methods. EditEval is a single-step editing benchmark consisting of original images, source captions, target captions, and editing prompts. The original images have resolutions ranging from 1863$\times$1863 to 8742$\times$8742. EditEval covers semantic editing (object addition, object removal, object replacement, and background change), stylistic editing (style change and texture change), and structural editing tasks (action change). Due to the extremely high resolutions of images, many editing tasks in EditEval require challenging fine-grained edits. An example is removing a tiny insect from a bird’s beak, which requires editing fewer than 1% of all pixels. EditEval allows a fair comparison of all methods, as none of them were trained on a dataset with the same distribution as EditEval.
Baseline Methods.
The baseline methods include InstructPix2Pix (IP2P) (Brooks et al., 2023), IP2P trained on MagicBrush (Zhang et al., 2024), and UltraEdit (Zhao et al., 2024). For these methods, we use the original editing prompts from EditEval. The inference hyperparameters are set to recommended values (Appendix A). For UltraEdit, we use the mask-based checkpoint and the same binary masks (with context dots) as in our workflow. Hence, UltraEdit similarly benefits from the extra user input, ensuring a fair comparison between our workflow and this strong baseline.
Evaluation Metrics.
We use the CLIP (Radford et al., 2021) text-image direction similarity (CLIPdir) and CLIP output similarity (CLIPout) metrics from the Emu Edit benchmark (Sheynin et al., 2024). These two metrics are suitable for reference-image-free evaluation. We do not calculate the L1 distance between edited and original images, as the L1 distance does not monotonically increase with the editing quality (higher distance can simutaneously indicate better global editing performance and worse local editing performance). In EditEval, we exclude the style change task, as the task is better handled by style LoRAs or specialized checkpoints (e.g., Flux.1 [dev] Redux [^9] https://huggingface.co/black-forest-labs/FLUX.1-Redux-dev). After excluding 24 images, there are $n=126$ images remaining.

### v2

@@ Abstract
Prompt-based models have demonstrated impressive prompt-following capability at image editing tasks. However, the models still struggle with following detailed editing prompts or performing local edits. Specifically, global image quality often deteriorates immediately after a single editing step. To address these challenges, we introduce SPICE, a training-free workflow that accepts arbitrary resolutions and aspect ratios, accurately follows user requirements, and consistently improves image quality during more than 100 editing steps, while keeping the unedited regions intact. By synergizing the strengths of a base diffusion model and a Canny edge ControlNet model, SPICE robustly handles free-form editing instructions from the user. On a challenging realistic image-editing dataset, SPICE quantitatively outperforms state-of-the-art baselines and is consistently preferred by human annotators. We release the workflow implementation for popular diffusion model Web UIs to support further research and artistic exploration. [^1] https://github.com/kenantang/spice
@@ 1 Introduction
Image editing is the task of changing the content of an image according to a user’s requirements. An example is adding an apple to a specific location on an image (Figure˜1). [^2] All figures are provided in high resolution. Readers are encouraged to zoom in to examine the details. A powerful image editing tool is vital for many applications from creative design to scientific research, including photo editing [17], video editing [6], data augmentation [9], and benchmark construction [26].
@@
[…]
@@
Existing vision-language models have achieved initial success on image editing [3, 29]. Such models take in an original image and the user’s editing prompt as the input and output the edited image, sometimes with a binary mask as additional input [25, 32]. However, for advanced artistic purposes that require more than one editing step, existing methods are disqualified by the following limitations. First, pixels outside the mask deteriorate after editing. Second, the user cannot specify the precise size and location of an added object by the mask. Third, models struggle with unusual editing tasks, such as adding a backpack to a bench facing away from the viewer (Figure˜3). Taken together, these limitations prohibit iterative editing, as image degradation inevitably accumulates. [^3] https://www.reddit.com/r/ChatGPT/comments/1kbj71z/i_tried_the_create_the_exact_replica_of_this/
To overcome these limitations, we propose SPICE, a training-free workflow consisting of 3 steps, namely mask generation, color and edge hint generation, and two-stage denoising (Figure˜2).
@@
First, in the mask generation step, the user provides a mask to define the editing region and the context size, localizing the modifications while using essential contextual information from the original image. SPICE strictly constrains the editing region to the user-provided mask, preventing deterioration.
Second, by accepting color and edge hints as a hinted image, SPICE allows the user to provide arbitrarily detailed or simplified image-space information to the model. This resolves the issue that textual prompts cannot specify precise sizes and locations [3, 29].
@@
Finally, SPICE uses a two-stage denoising process, where a Canny edge ControlNet model [30] integrates image-space hints in the early denoising steps, and a base diffusion model refines and diversifies the output in the later steps. By synergizing the complementary strengths of both models, the two-stage denoising process enables precise and customizable editing, in which the model faithfully follows the user’s requirements on properties of the edited object. Beyond image editing, the superior faithfulness also helps users to overcome fundamental limitations of AI-generated artwork, such as predominantly portraying a single character or repeatedly erring on details.
@@
By overcoming the limitations, SPICE consistently achieves success in iterative editing tasks of more than 100 editing steps (Section˜B.3). For the first time, our workflow enables scaling test-time compute [24] in image generation, where the user can decide how each unit of additional compute time contributes to the final image quality. This user-friendly design distinguishes our workflow from contemporary test-time scaling approaches that rely on automatic search algorithms [16, 23].
@@
Despite the strong capability, our workflow can be easily integrated into all popular diffusion model Web UIs, such as Stable Diffusion Web UI Automatic1111, ComfyUI, and Stable Diffusion Web UI Forge. Unlike existing tools such as ADetailer [^4] https://github.com/Bing-su/adetailer or Regional Prompter, [^5] https://github.com/hako-mikan/sd-webui-regional-prompter which offer an overwhelming number of hyperparameters that can be difficult to navigate, SPICE provides a much smaller set of hyperparameters for ease-of-use. Moreover, SPICE introduces minimal computational overhead in each editing step, making editing as efficient as generating an image from text with the same model and sampling hyperparameters. This extremely low overhead allows SPICE to run on consumer GPUs (e.g., a single NVIDIA GeForce RTX 4090), unlocking iterative editing for more users. Combining strong capabilities with low implementation and computational cost, our workflow provides a powerful yet accessible tool for researchers and non-researchers alike.
@@
Figure 2: By sketching a binary mask with context dots and a color & edge hint, users can effortlessly achieve realistic edits with SPICE. Subfigure (a) shows the overview of our workflow, while Subfigures (b) and (c) show the internal steps. In this example, the user requires a sunhat to be added next to the woman. First, the user sketches both a mask with context dots and a hinted image containing color and edge hints. The mask is automatically blurred after being sketched. Then, during the two-stage denoising step, the Canny and base models perform the early and late denoising steps, respectively. See Figure˜18 for more examples of masks and hints.
@@ 2 Methods
SPICE is based on inpainting [15], an operation that uses a diffusion model to replace a masked region on an image, conditioned on a prompt that describes the new image. Note that this prompt differs from the editing prompt, which uses a verb (e.g., add or remove) to describe the editing operation [29]. For example, when the editing prompt is “add an apple in the fridge”, the description prompt will be “an apple in a fridge”. In this section, we introduce the three key steps of our workflow, namely mask generation, color and edge hint generation and two-stage denoising (Figure˜2).
@@ 2.1 Mask Generation
[…]
@@
To address these issues, we introduce context dots, a pair of dots at opposite corners of the desired bounding box. Context dots ensure that the extended bounding box includes sufficient context for generation. The user directly adds context dots to the original mask, resulting in a context mask $M_{\text{context}}\in\{0,1\}^{H\times W}$, as shown in Figure 2(b). This simple enhancement offers three key advantages: (1) the user can exclude image areas that interfere with the inpainting process, (2) the user can specify a resolution between that of the inpainted region and that of the full image, balancing local details and global context (Section˜B.4), and (3) context dots minimally affect surrounding pixels, limiting changes to only the desired editing region.
@@
On inpainted images, there is usually an unwanted sharp boundary between the inpainted region and the remaining parts of the image. In more severe cases, an inpainted object can be incomplete. The reason is that even with a context, diffusion models are not robust enough to generate pixels that seamlessly blend with existing ones. To mitigate these artifacts, we adopt Differential Diffusion [12], a method that allows the diffusion model to be conditioned by continuous mask values in [0, 1] during generation. We provide the continuous values as a soft mask $M_{\text{soft}}\in[0,1]^{H\times W}$ by applying a simple Gaussian blur to $M_{\text{context}}$ with a kernel size of a few pixels, as shown in Figure 2(b). This procedure, together with thresholds for blending the original and inpainted image, is named as Soft Inpainting in popular Web UIs. We use Soft Inpainting when it is available in a Web UI or implement our own simplified version (without thresholds) when it is not available.
@@ 2.2 Color and Edge Hint Generation
In many existing image editing systems [5, 27, 10], users typically specify desired content through a textual prompt. However, words alone cannot easily describe certain details, such as asymmetric apparel designs or intricate color patterns. Instead, these subtle details can be effectively conveyed with an additional visual hint in the image space. To this end, SPICE allows a user to first create a rough sketch or color layout using standard editing software (e.g., Krita or Adobe Photoshop). The resultant image, denoted $I_{\text{hinted}}\in[0,1]^{H\times W\times C}$, then replaces the original image as the inpainting input. To incorporate information from this hinted image, a denoising strength hyperparameter in [0, 1] specifies how much the original pixels on the image should be changed. [^7] The hyperparameter does not work as a simple linear blending coefficient at the post-processing stage, and its exact implementation differs for various algorithms. To avoid ambiguity, we refrain from providing a general equation here. Interested readers can refer to the source code of popular Web UIs for the equations. At a moderate denoising strength (e.g., 0.5 to 0.7), the inpainting model produces pixels that remain close to the user’s sketched hints, preserving intended colors, shapes, or patterns. Meanwhile, the pixels still diverges enough from the sketch, forming realistic objects in the end.
Note that the user can provide any form of color and edge hints, including sketches, reference images pasted in a collage style, or even another region of the original image (such as using the Clone Stamp Tool in PhotoShop). This flexibility is an improvement over existing methods, such as MagicQuill [14] that uses downsampled $\times32$ color blocks for guidance and thus loses high-frequency details. Hence, SPICE allows a user to perform image editing from any point on the human-model collaboration spectrum. At one end, the user fully edit the image by drawing out every detail, without the help of diffusion models. At the other end, the user fully delegates the model to edit the image. Usually, the user can get decent editing results by staying on the end where the user input is minimal (Section˜B.1). We will also discuss how a small set of hyperparameters enable the user to move freely along this spectrum (Section˜B.4).
@@ 2.3 Two-Stage Denoising
Image editing methods [15] typically rely on a single diffusion model to perform all denoising steps when generating an image from latent noise. However, different diffusion models have complementary strengths. On the one hand, a general-purpose text-to-image base model, such as Flux.1 [dev], [^8] https://huggingface.co/black-forest-labs/FLUX.1-dev excels at generating rich variations in its output but can only be conditioned on textual prompts. On the other hand, Flux.1 [dev] Canny, [^9] https://huggingface.co/black-forest-labs/FLUX.1-Canny-dev a model derived from Flux.1 [dev] that contains a Canny edge ControlNet (i.e., a Canny model), can be conditioned on Canny edge information [30] but sacrifices variability.
@@
Empirically, we find that for a wide range of image editing tasks, Canny Edge ControlNet models outperforms other ControlNet variants (e.g., depth or pose) in the first denoising stage. Furthermore, Canny models are more widely available than other ControlNet models. Therefore, we only use Canny models but not other ControlNet models in our workflow.
@@ 3 Experiments
We comprehensively evaluate SPICE on challenging editing tasks. Specifically, we stress test the three features of SPICE, namely precise (Section˜B.2), iterative (Section˜B.3), and customizable editing (Section˜B.4). We also provide ablation studies to justify each step in SPICE. Due to page limitations, full details and more results are deferred to Appendix˜B. In this section, we highlight the results that demonstrate the unique strengths of SPICE over state-of-the-art baselines.
3.1 Benchmark Results
We evaluate SPICE against baselines of representative open and proprietary models. The results show that SPICE qualitatively outperforms both open (Figure˜3) and proprietary models (Appendix˜E), with a clear advantage in its prompt-following capability.
3.2 Precise, Iterative, and Customizable Editing
Prompt-based image generation or editing models frequently fail to fulfill their promise that users can create content according to their own will. Users are frustrated for two main reasons.
First, prompts are hard to design and cannot precisely instruct the model to produce complex and unusual outputs (Figure˜4(a)). A user assumes that the model can follow the prompt, but the model always fails to interpret the prompt as humans do, even after the user repeatedly corrects the model.
Original IP2P MB UE Ours Original IP2P MB UE Ours
@@ 3.2 Precise, Iterative, and Customizable Editing
Figure 3: Our workflow outperforms baseline methods in 6 editing categories from EditEval. Each group of five images shows an example from an editing task. Each group starts from original image, followed by IP2P, MagicBrush (MB), UltraEdit (UE), and our results.
Round 1 Round 1
Refer to caption Refer to caption
Round 2 Round 2
Refer to caption Refer to caption
Round 3 Round 3
Refer to caption Refer to caption
(a) DALL·E 3
(b) GPT-4o
(c) Ours
Figure 4: SPICE can generate content that DALL·E 3 and GPT-4o cannot. Two examples are a rabbit with 4 ears and a violin without a bridge. For individual objects, DALL·E 3 fails even after the user asks the model to edit the errors multiple times. For combined objects, GPT-4o cannot generate all objects correctly at once. However, SPICE can reliably generate these challenging objects.
Second, the model’s randomness makes complex images costly—and even impossible—to generate. For example, if an image must include 10 objects and each is generated correctly only 50% of the time, the chance of getting all 10 right drops below 0.1%. In practice, object-level correctness rate is even lower and scene complexity higher, so achieving perfect results becomes impossible (Figure˜4(b)).
With the high customizability of SPICE, instead of desperately engineering prompts, users can effortlessly generate compositions and objects that are impossible for DALL·E 3 and GPT-4o (Figure˜4, see full prompt in Section˜B.4). [^10] Figure 4(c) has been accepted to CVPR AI Art Gallery 2025. https://thecvf-art.com/project/compositionality-and-parts/ Hence, SPICE creates a new paradigm for human-model collaboration, where users can fully realize their creative vision, without ever needing to compromise to the models’ lack of prompt compliance. More instructions and examples of iterative and customizable image generation and editing can be found in Appendix˜A, Appendix˜J, and Appendix˜L.
4 Related Work
Diffusion model-based image editing techniques can be categorized based on the type of inputs used to guide image generation [5, 27, 10]. Most existing methods rely primarily on textual prompts [3, 7], while many also incorporate masks to specify the editing region [25, 32]. Due to the inherent imprecision of both prompts and masks, some methods integrate additional inputs to provide finer control over the generation process [28, 22, 13]. Among these, MagicQuill [14] is most closely related to our workflow. MagicQuill similarly enables user-guided editing using color and edge information. However, its color guidance is constrained to low-resolution $\times32$ blocks, regardless of image size, prohibiting fine-grained edits. SPICE overcomes this limitation of MagicQuill by accepting full-resolution color and edge hints, effectively interpreting the provided guidance and leading to precise edits. More discussion to compare SPICE and MagicQuill can be found in Appendix˜M.
5 Conclusion
Existing prompt-based image editing models fail in performing local edits, working under different resolutions, following user instructions, and maintaining image quality during multiple editing steps. We propose SPICE, a training-free workflow that addresses all these challenges. We release the workflow to facilitate future research and artistic exploration.
Acknowledgments
We thank Anthony Yang for his assistance in preparing the materials presented in Appendix˜J.
Appendix A Hints and Prompts of the Fridge Example
In Table˜1, we list the simple hints and prompts that we use to produce results in Figure˜1. We also explain how each hint is quickly and easily added to the image by using Photoshop. The explanation describes the difference between the hinted image at Step $i$ and the result image at Step i1i-1. The results demonstrate two advantages of SPICE. On the one hand, SPICE is highly robust to different types of hints. On the other hand, SPICE allows user to effortlessly achieve sophisticated editing results, without the need to optimize complicated hints or prompts. Sometimes, prompts do not even need to be changed over different steps, as SPICE can infer the user’s intent from the hints.
Table 1: To generate realistic edits, a user only needs to provide simple hints and prompts. For each step in Figure˜1, the image with hints (“Hinted”) and the description prompt are shown below. We also explain the hint for each step. The explanation is not part of the input to the model.
Step Hinted Result Description Prompt Hint Explanation
1 [Uncaptioned image] [Uncaptioned image] An open fridge with food in it. The contours of the fridge door and fridge shelves are roughly sketched using the hard round brush.
2 [Uncaptioned image] [Uncaptioned image] An open fridge with food in it. The region occupied by the removed bottle is recolored using the hard round brush and the Clone Stamp tool.
3 [Uncaptioned image] [Uncaptioned image] An open fridge with food in it. A single red circle is added using the hard round brush.
4 [Uncaptioned image] [Uncaptioned image] An open fridge with food in it. A beer. First, the coupe glass is removed using the Clone Stamp tool. Then, an image of Monster Energy is pasted at the same place.
5 [Uncaptioned image] [Uncaptioned image] An open fridge with food in it. A bowl with a word “suspicious” on it. Ten black dots corresponding to ten letters of “suspicious” are added on the bowl using the hard round brush.
6 [Uncaptioned image] [Uncaptioned image] An open fridge with food in it. A purple potato. The potato is selected by the Quick Selection tool. The hue of the selected region is changed by the Hue/Saturation tool.
7 [Uncaptioned image] [Uncaptioned image] An open fridge with food in it. From the image in Step 0, a thin region above the fridge door is selected using the Quick Selection tool and copied to the image.
8 [Uncaptioned image] [Uncaptioned image] An open fridge with food in it. From the image in Step 0, a thin region below the fridge door is selected using the Quick Selection tool and copied to the image.
9 [Uncaptioned image] [Uncaptioned image] An open fridge with food in it. No change is made.
Appendix B Overview of Results
Original IP2P MB UE Ours Original IP2P MB UE Ours
Figure 5: Our workflow outperforms baseline methods in 6 editing categories from EditEval. Each group of five images shows an example from an editing task. Each group starts from original image, followed by IP2P, MagicBrush (MB), UltraEdit (UE), and our results.
B.1 Benchmark Results